using a single hose janet can fill a pool in 7 hours. the same pool can be drained in 8 hours by opening a drainpipe. if janet forgets to close the drainpipe how long would it take her to fill the pool.
Answer by lwsshak3(11628)

using a single hose janet can fill a pool in 7 hours. the same pool can be drained in 8 hours by opening a drainpipe.
***
let x=amt of time it would take to fill the pool.
1/7=fill rate
1/8=drain rate
x/7-x/8=1
lcd:7*8=56
8x-7x=56
x=56
if janet forgets to close the drainpipe how long would it take her to fill the pool. 56 hrs